OpenStudy (anonymous):

The sales price if a sweater is 20% less than the original price. If the sale price was $48, what was the original price?

OpenStudy (anonymous):

48-(48*0.20) = 48 - 9.6 = 38.4

OpenStudy (anonymous):

I don't think this is right. The sale price of the sweater can't be less than the original price.

OpenStudy (anonymous):

original price/100% = 48 (sale price)/80% x/100=48/80 solve the proportion and you'll have your original price